### v3

@@ 1. Introduction
In Section 5, we introduce the dominant sets for the seeds of cluster algebras as an replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory and as an replacement of inversion sets in the study of right weak order on Weyl groups. Then we prove the results on oriented exchange graphs. The main results in this section are Theorems 5.7, 5.12 and Propositions 5.21, 5.30.
@@ 2. Preliminaries on cluster algebras
Throughout this paper, we fix a pair $(n,m)$ of integers with $m\geq n>0$. For any positive integer $r$, we denote $[1,r]:=\{1,\ldots,r\}$. The matrices in this paper are always integer matrices.
@@ 2.1. Mutation matrices and their mutations
where $[a]_{+}:=\max\{a,0\}$ for any $a\in\mathbb{R}$.
The following statements are well-known, c.f., [fz_2002, Proposition 4.5], [bfz_2005, Lemma 3.2].
@@
  • The rank of $\widetilde{B}^{\prime}=\mu_{k}(\widetilde{B})$ is equal to the rank of $\widetilde{B}$.
As shown in [bfz_2005, (3.2)], the matrix $\widetilde{B}^{\prime}=\mu_{k}(\widetilde{B})$ can be written as
@@ 2.2. Compatible pairs and their mutations
Notice that the diagonal matrix $S$ is uniquely determined by the compatible pair $(\widetilde{B},\Lambda)$. The matrix $\Lambda$ in a compatible pair $(\widetilde{B},\Lambda)$ is called a quantization matrix.
Proposition 2.2.
[bz-2005, Proposition 3.3] Let $(\widetilde{B},\Lambda)$ be a compatible pair of type $S$ and $B$ the top $n\times n$ submatrix of $\widetilde{B}$. Then the matrix $\widetilde{B}$ has the full rank $n$ and $SB=\widetilde{B}^{T}\Lambda\widetilde{B}$. In particular, $B$ is skew-symmetrizable.
@@
Let $(\widetilde{B},\Lambda)$ be a compatible pair of type $S$ and $k\in[1,n]$. We set $\widetilde{B}^{\prime}:=\mu_{k}(\widetilde{B})=E_{k,\varepsilon}^{\widetilde{B}}\widetilde{B}F_{k,\varepsilon}^{\widetilde{B}}$ and
$\Lambda^{\prime}:=(E_{k,\varepsilon}^{\widetilde{B}})^{T}\Lambda E_{k,\varepsilon}^{\widetilde{B}}.$
Proposition 2.3.
[bz-2005, Proposition 3.4] (i) The matrix $\Lambda^{\prime}$ is independent of the choice of $\varepsilon\in\{\pm\}$.
(ii) The pair $(\widetilde{B}^{\prime},\Lambda^{\prime})$ is still a compatible pair of type $S$.
@@
The following proposition will be not used in this paper, but it is very useful to construct many examples of compatible pairs from a given skew-symmetrizable matrix.
Proposition 2.4.
[zelevinsky-2005, Example 0.5] Let $B$ be an $n\times n$ skew-symmetrizable integer matrix and set $\widetilde{B}=\begin{bmatrix}B\\ I_{n}\end{bmatrix}$. Then the skew-symmetric matrices $\Lambda$ compatible with $\widetilde{B}$ in the sense of (2.4) are those of the form
@@ 2.3. $Y$-pattern and cluster pattern
A $Y$-seed of rank $n$ in $\mathbb{F}$ is a pair $({\bf y},\widehat{B})$, where
@@
The variables $y_{1},\ldots,y_{m}$ are called the $y$-variables of $({\bf y},\widehat{B})$.
Let $({\bf y},\widehat{B})$ be a $Y$-seed of rank $n$ in $\mathbb{F}$. The mutation of $({\bf y},\widehat{B})$ in direction $k\in[1,n]$ is the pair $({\bf y}^{\prime},\widehat{B}^{\prime}):=\mu_{k}({\bf y},\widehat{B})$ given as follows:
@@
One can check that the new pair $({\bf y}^{\prime},\widehat{B}^{\prime})=\mu_{k}({\bf y},\widehat{B})$ is still a $Y$-seed of rank $n$ and $({\bf y},\widehat{B})=\mu_{k}({\bf y}^{\prime},\widehat{B}^{\prime})$.
A cluster seed or simply a seed of rank $n$ in $\mathbb{F}$ is a pair $({\bf x},\widetilde{B})$, where
@@
In this case, the tuple ${\bf x}$ is called the cluster of $({\bf x},\widetilde{B})$. Elements in ${\bf x}$ are called cluster variables. More precisely, we call $x_{1},\ldots,x_{n}$ unfrozen cluster variables and $x_{n+1},\ldots,x_{m}$ frozen cluster variables. The matrices $B$ and $P$ are respectively called the exchange matrix, coefficient matrix of $({\bf x},\widetilde{B})$. We denote by $\widehat{\bf y}:=(\hat{y}_{1},\ldots,\hat{y}_{n})$, where $\hat{y}_{k}:=\prod_{j=1}^{m}x_{j}^{b_{jk}}$. The variables $\hat{y}_{1},\ldots,\hat{y}_{n}$ are called $\hat{y}$-variables associated to $({\bf x},\widetilde{B})$.
Let $({\bf x},\widetilde{B})$ be a seed of rank $n$ in $\mathbb{F}$. The mutation of $({\bf x},\widetilde{B})$ in direction $k\in[1,n]$ is the pair $({\bf x}^{\prime},\widetilde{B}^{\prime})=\mu_{k}({\bf x},\widetilde{B})$ given by $\widetilde{B}^{\prime}=\mu_{k}(\widetilde{B})$ and
@@
One can check that the new pair $({\bf x}^{\prime},\widetilde{B}^{\prime})=\mu_{k}({\bf x},\widetilde{B})$ is still a seed of rank $n$ and $({\bf x},\widetilde{B})=\mu_{k}({\bf x}^{\prime},\widetilde{B}^{\prime})$.
@@
Let $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$ be a cluster pattern, $\mathcal{S}_{Y}=\{({\bf y}_{t},\widehat{B}_{t})\mid t\in\mathbb{T}_{n}\}$ a $Y$-pattern, and let $\Lambda=\{\Lambda_{t}\mid t\in\mathbb{T}_{n}\}$ be a collection of $m\times m$ skew-symmetric integer matrices indexed by the vertices in $\mathbb{T}_{n}$.
@@
    The triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\Lambda)$ is called a Langlands-Poisson triple, if it satisfies that
    • $(\mathcal{S}_{X},\mathcal{S}_{Y})$ forms a Langlands dual pair;
    • $\{(\widetilde{B}_{t},\Lambda_{t})\mid t\in\mathbb{T}_{n}\}$ forms a collection of compatible pairs and $(\widetilde{B}_{t^{\prime}},\Lambda_{t^{\prime}})=\mu_{k}(\widetilde{B}_{t},\Lambda_{t})$ whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$.
The above definition is purely for convenience of this paper. The following result can be checked easily.
@@
[…]
$\{x_{i;t_{0}},x_{j;t_{0}}\}:=\lambda_{ij;t_{0}}\cdot x_{i;t_{0}}x_{j;t_{0}}.$
@@
    $\mathcal{A}:=\mathbb{Z}[x_{1;t},\ldots,x_{n;t},x_{n+1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}\mid t\in\mathbb{T}_{n}].$
@@
    $\mathcal{U}:=\bigcap\limits_{t\in\mathbb{T}_{n}}\mathcal{L}(t),$
    where $\mathcal{L}(t):=\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$.
Theorem 2.9.
[fz_2002]*Laurent phenomenon For any cluster variable $z$ and any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{A}$, we have $z\in\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{n;t}^{\pm 1},x_{n+1;t},\ldots,x_{m;t}]$. In particular, $\mathcal{A}$ is contained in $\mathcal{U}$.
By the Laurent phenomenon, any cluster variable $z$ can be uniquely written as
$z=\frac{N(x_{1;t},\ldots,x_{n;t})}{x_{1;t}^{d_{1}^{t}}\ldots x_{n;t}^{d_{n}^{t}}},$
where $N$ is a polynomial in $x_{1;t},\ldots,x_{n;t}$ with coefficents in $\mathbb{Z}[x_{n+1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$ with $x_{i;t}\nmid N$ for $i\in[1,n]$, and each $d_{i}^{t}$ is an integer. The vector ${\bf d}_{z}^{t}:=(d_{1}^{t},\ldots,d_{n}^{t})\in\mathbb{Z}^{n}$ is called the $d$-vector of $z$ with respect to seed at the vertex $t$.
@@
Take $\widetilde{B}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ and ${\bf x}=(x_{1},x_{2})$. One can check that the cluster algebra $\mathcal{A}$ defined by the initial seed $({\bf x},\widetilde{B})$ has only five cluster variables:
$x_{1},\;x_{2},\;x_{3}:=\frac{x_{2}+1}{x_{1}},\;x_{4}:=\frac{x_{1}+x_{2}+1}{x_{1}x_{2}},\;x_{5}:=\frac{x_{1}+1}{x_{2}}.$
It can be seen that all the cluster variables are in $\mathbb{Z}[x_{1}^{\pm 1},x_{2}^{\pm 1}]$. The $d$-vectors with respect to the initial seed $({\bf x},\widetilde{B})$ are as follows:
${\bf d}_{x_{1}}=\begin{bmatrix}-1\\ 0\end{bmatrix},\;\;{\bf d}_{x_{2}}=\begin{bmatrix}0\\ -1\end{bmatrix},\;\;{\bf d}_{x_{3}}=\begin{bmatrix}1\\ 0\end{bmatrix},\;\;{\bf d}_{x_{4}}=\begin{bmatrix}1\\ 1\end{bmatrix},\;\;{\bf d}_{x_{5}}=\begin{bmatrix}0\\ 1\end{bmatrix}.$
@@ 3. Tropical points, pointed elements and good elements
3.1. Universal semi-field and tropical points
Recall that $(\mathbb{P},\cdot,\oplus)$ is called a semi-field if $(\mathbb{P},\cdot)$ is an abelian multiplicative group endowed with a binary operation of auxiliary addition $\oplus$ which is commutative, associative and satisfies that the multiplication distributes over the auxiliary addition. For example,
$\mathbb{Z}^{\rm min}:=(\mathbb{Z},+,\min\{-,-\})\;\;\;\text{and}\;\;\;\mathbb{Z}^{\rm max}:=(\mathbb{Z},+,\max\{-,-\})$
are semi-field and they are called tropical semi-field.
Let $\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ be the set of all non-zero rational functions in $u_{1},\ldots,u_{m}$ that have subtraction free expressions. The set $\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ is a semifield with respect to the usual operations of multiplication and addition. It is called an universal semi-field.
From now on, we fix $\mathbb{F}:=\mathbb{Q}(u_{1},\ldots,u_{m})$ and $\mathbb{F}_{>0}:=\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$. For any semi-field $\mathbb{P}$ and ${\bf p}=(p_{1},\ldots,p_{m})^{T}\in\mathbb{P}^{m}$, there exists a unique semi-field homomorphism
@@ 3.1. Universal semi-field and tropical points
induced by $u_{i}\mapsto p_{i}$ for any $i$. The map $\pi:{\bf p}\mapsto\pi_{\bf p}$ induces a bijection from $\mathbb{P}^{m}$ to the set $\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{P})$ of semifield homomorphisms from $\mathbb{F}_{>0}$ to $\mathbb{P}$.
@@
Let $\mathcal{C}=\{{\bf u}_{t}\mid t\in\mathbb{T}\}$ be a collection of charts in $\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ indexed by any set $\mathbb{T}$. Then the following statements hold.
@@
    For any chart ${\bf u}_{t}=(u_{1;t},\ldots,u_{m;t})$, denote by $u_{i}({\bf u}_{t})$ the expression of $u_{i}$ as an element in $\mathbb{F}(t)_{>0}:=\mathbb{Q}_{\rm sf}(u_{1;t},\ldots,u_{m;t})$. The map $\varphi_{t}:u_{i}\mapsto u_{i}({\bf u}_{t})$ induces an isomorphism of semifield from $\mathbb{F}_{>0}$ to $\mathbb{F}(t)_{>0}$.
@@
    ${\bf q}^{t}(\beta):=(\beta(u_{1;t}),\ldots,\beta(u_{m;t}))^{T}\in\mathbb{Z}^{m}.$
@@
Proof.
Obviously. ∎
Keep the setting in the above proposition. We call the vector
@@
  • If we work on a $Y$-pattern $\mathcal{S}_{Y}=\{({\bf y}_{t},\widehat{B}_{t})\mid t\in\mathbb{T}_{n}\}$, we take $\mathbb{F}_{>0}$ to be the universal semifield generated by the initial $y$-variables $y_{1;t_{0}},\ldots,y_{m;t_{0}}$. Then $\mathcal{C}:=\{{\bf y}_{t}\mid t\in\mathbb{T}_{n}\}$ is a collection of charts in $\mathbb{F}_{>0}$. We usually denote by
@@
  • If we work on a cluster pattern $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$, we take $\mathbb{F}_{>0}$ to be the universal semifield generated by the initial cluster variables $x_{1;t_{0}},\ldots,x_{m;t_{0}}$. Then $\mathcal{C}:=\{{\bf x}_{t}\mid t\in\mathbb{T}_{n}\}$ is a collection of charts in $\mathbb{F}_{>0}$. We usually denote by
@@
    A tropical point $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ associated to the $Y$-pattern $\mathcal{S}_{Y}$ is an assignment of a (column) vector ${\bf g}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ in $\mathbb{Z}^{m}$ to each vertex $t$ of $\mathbb{T}_{n}$ such that
@@
    A tropical point $[{\bf a}]=\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ associated to the cluster pattern $\mathcal{S}_{X}$ is an assignment of a (column) vector ${\bf a}^{t}=(a_{1}^{t},\ldots,a_{m}^{t})^{T}$ in $\mathbb{Z}^{m}$ to each vertex $t$ of $\mathbb{T}_{n}$ such that
@@
    Each tropical point $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ corresponds to a unique semifield homomorphism $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ such that
@@
    Each tropical point $[{\bf a}]=\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ corresponds to a unique semifield homomorphism $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ such that

### v4

@@ 1.6. Organization
In Section 7, we introduce the dominant sets for the seeds of cluster algebras as a replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory. Then we prove the results on oriented exchange graphs (see Theorem 7.8 and Theorem 7.13).
@@ 2. Preliminaries on cluster algebras
Throughout this paper, we fix a pair $(n,m)$ of integers with $m\geq n>0$. For any positive integer $r$, we denote $1,r]\coloneqq\{1,\ldots,r\}$. The matrices in this paper are always integer matrices.
@@ 2.1. Mutation matrices and their mutations
where $a]_{+}\coloneqq\max\{a,0\}$ for any $a\in\mathbb{R}$.
The following statements are well-known, cf. [fz_2002, Proposition 4.5], [bfz_2005, Lemma 3.2].
@@
[…]
@@ 2.2. Compatible pairs and their mutations
Notice that the diagonal matrix $S$ is uniquely determined by the compatible pair $(\widetilde{B},\Lambda)$. The matrix $\Lambda$ in a compatible pair $(\widetilde{B},\Lambda)$ is called a Poisson coefficient matrix.
Proposition 2.2 ([bz-2005, Proposition 3.3]).
Let $(\widetilde{B},\Lambda)$ be a compatible pair of type $S$ and $B$ the top $n\times n$ submatrix of $\widetilde{B}$. Then the matrix $\widetilde{B}$ has the full rank $n$ and $SB=\widetilde{B}^{T}\Lambda\widetilde{B}$. In particular, $B$ is skew-symmetrizable.
@@
Let $(\widetilde{B},\Lambda)$ be a compatible pair of type $S$ and $k\in[1,n]$. We set $\widetilde{B}^{\prime}\coloneqq\mu_{k}(\widetilde{B})=E_{k,\varepsilon}^{\widetilde{B}}\widetilde{B}F_{k,\varepsilon}^{\widetilde{B}}$ and
$\Lambda^{\prime}\coloneqq(E_{k,\varepsilon}^{\widetilde{B}})^{T}\Lambda E_{k,\varepsilon}^{\widetilde{B}}=(\lambda_{ij}^{\prime}).$
We have
(2.5) $\lambda_{ij}^{\prime}=\begin{cases}-\lambda_{ik}+\sum_{l=1}^{m}[-\varepsilon b_{lk}]_{+}\lambda_{il},&\text{if }j=k;\\ -\lambda_{kj}+\sum_{l=1}^{m}[-\varepsilon b_{lk}]_{+}\lambda_{lj},&\text{if }i=k;\\ \lambda_{ij},&\text{otherwise}.\end{cases}$
Proposition 2.3 ([bz-2005, Proposition 3.4]).
  • (i)
    The matrix $\Lambda^{\prime}$ is independent of the choice of $\varepsilon\in\{\pm\}$.
  • (ii)
    The pair $(\widetilde{B}^{\prime},\Lambda^{\prime})$ is still a compatible pair of type $S$.
@@
The following proposition is very useful to construct many examples of compatible pairs from a given skew-symmetrizable matrix.
Proposition 2.4 ([zelevinsky-2005, Example 0.5]).
Let $B$ be an $n\times n$ skew-symmetrizable integer matrix and set $\widetilde{B}=\begin{bmatrix}B\\ I_{n}\end{bmatrix}$. Then the skew-symmetric matrices $\Lambda$ compatible with $\widetilde{B}$ in the sense of (2.4) are those of the form
@@ 2.3. $Y$-pattern and cluster pattern
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Let $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$ be a cluster pattern, $\mathcal{S}_{Y}=\{({\bf y}_{t},\widehat{B}_{t})\mid t\in\mathbb{T}_{n}\}$ a $Y$-pattern, and let $\mathsf{\Lambda}=\{\Lambda_{t}\mid t\in\mathbb{T}_{n}\}$ be a collection of $m\times m$ skew-symmetric integer matrices indexed by the vertices in $\mathbb{T}_{n}$.
@@
    The pair $(\mathcal{S}_{X},\mathsf{\Lambda})$ is called a $\mathsf{\Lambda}$-cluster pattern, if $\{(\widetilde{B}_{t},\Lambda_{t})\mid t\in\mathbb{T}_{n}\}$ forms a collection of compatible pairs and $(\widetilde{B}_{t^{\prime}},\Lambda_{t^{\prime}})=\mu_{k}(\widetilde{B}_{t},\Lambda_{t})$ whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$. In this case, we call the triple $({\bf x}_{t},\widetilde{B}_{t},\Lambda_{t})$ a $\mathsf{\Lambda}$-seed.
  • (iii)
    The triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$ is called a Langlands-Poisson triple, if $(\mathcal{S}_{X},\mathsf{\Lambda})$ is a $\mathsf{\Lambda}$-cluster pattern and $(\mathcal{S}_{X},\mathcal{S}_{Y})$ is a Langlands dual pair.
The above definition is purely for convenience of this paper. Notice that each cluster pattern has a unique Langlands dual $Y$-pattern up to isomorphism. The following result can be checked easily.
@@
Given a $\mathsf{\Lambda}$-cluster pattern $(\mathcal{S}_{X},\mathsf{\Lambda})$, we can define a Poisson bracket $\{-,-\}$ on the ambient field $\mathbb{F}$ using the $m\times m$ skew-symmetric matrix $\Lambda_{t_{0}}=(\lambda_{ij;t_{0}})$ at the rooted vertex $t_{0}$:
$\{x_{i;t_{0}},x_{j;t_{0}}\}\coloneqq\lambda_{ij;t_{0}}\cdot x_{i;t_{0}}x_{j;t_{0}}.$
@@
    $\mathcal{A}\coloneqq\mathbb{Z}[x_{1;t},\ldots,x_{n;t},x_{n+1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}\mid t\in\mathbb{T}_{n}].$
    Denote by $\mathcal{A}^{+}\coloneqq\mathbb{Z}[x_{1;t},\ldots,x_{m;t}\mid t\in\mathbb{T}_{n}]\subseteq\mathcal{A}$, the version of cluster algebra with frozen variables non inverted.
@@
    $\mathcal{U}\coloneqq\bigcap\limits_{t\in\mathbb{T}_{n}}\mathcal{L}(t),$
    where $\mathcal{L}(t)\coloneqq\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$.
  • (iii)
    If $(\mathcal{S}_{X},\mathsf{\Lambda})$ is a $\mathsf{\Lambda}$-cluster pattern, then the corresponding (upper) cluster algebra $\mathcal{A}$, $\mathcal{A}^{+}$ and $\mathcal{U}$ are endowed with an extra data $\mathsf{\Lambda}=\{\Lambda_{t}\mid t\in\mathbb{T}_{n}\}$. In this case, we call $\mathcal{A}$ or $\mathcal{A}^{+}$ a $\mathsf{\Lambda}$-cluster algebra and call $\mathcal{U}$ a $\mathsf{\Lambda}$-upper cluster algebra.
  • (iv)
    If $m=n$, i.e., there are no frozen variables, the corresponding (upper) cluster algebra is said to be with trivial coefficients.
Notice that the mutation matrices of $\mathsf{\Lambda}$-(upper) cluster algebras always have full rank, by Proposition 2.2.
Theorem 2.9 ([fz_2002]*Laurent phenomenon).
For any cluster variable $z$ and any seed $({\bf x}_{t},\widetilde{B}_{t})$ of a cluster algebra $\mathcal{A}$, we have
$z\in\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{n;t}^{\pm 1},x_{n+1;t},\ldots,x_{m;t}].$
In particular, $\mathcal{A}$ is contained in $\mathcal{U}$.
@@
[…]
$x_{1},\;x_{2},\;x_{3}\coloneqq\frac{x_{2}+1}{x_{1}},\;x_{4}\coloneqq\frac{x_{1}+x_{2}+1}{x_{1}x_{2}},\;x_{5}\coloneqq\frac{x_{1}+1}{x_{2}}.$
It is clear that all the cluster variables are contained in $\mathbb{Z}[x_{1}^{\pm 1},x_{2}^{\pm 1}]$.
@@ 3. Tropical points, pointed elements and good elements
3.1. Universal semifield and tropical points
Recall that $(\mathbb{P},\cdot,\oplus)$ is called a semifield if $(\mathbb{P},\cdot)$ is an abelian multiplicative group endowed with a binary operation of auxiliary addition $\oplus$ which is commutative, associative and satisfies that the multiplication distributes over the auxiliary addition. For example,
$\mathbb{Z}^{\rm min}\coloneqq(\mathbb{Z},+,\min\{-,-\})\;\;\;\text{and}\;\;\;\mathbb{Z}^{\rm max}\coloneqq(\mathbb{Z},+,\max\{-,-\})$
are semifield and they are called tropical semifield.
Let $\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ be the set of all non-zero rational functions in $u_{1},\ldots,u_{m}$ that have subtraction free expressions. The set $\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ is a semifield with respect to the usual operations of multiplication and addition. It is called an universal semifield.
From now on, we fix $\mathbb{F}\coloneqq\mathbb{Q}(u_{1},\ldots,u_{m})$ and $\mathbb{F}_{>0}\coloneqq\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$. For any semifield $\mathbb{P}$ and ${\bf p}=(p_{1},\ldots,p_{m})^{T}\in\mathbb{P}^{m}$, there exists a unique semifield homomorphism
@@ 3.1. Universal semifield and tropical points
induced by $u_{i}\mapsto p_{i}$ for any $i$. The map $\pi\colon{\bf p}\mapsto\pi_{\bf p}$ induces a bijection from $\mathbb{P}^{m}$ to the set $\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{P})$ of semifield homomorphisms from $\mathbb{F}_{>0}$ to $\mathbb{P}$.
@@
Let $\mathcal{C}=\{{\bf u}_{t}\mid t\in\mathbb{T}\}$ be a collection of charts in $\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ indexed by a set $\mathbb{T}$. Then the following statements hold.
@@
    For any chart ${\bf u}_{t}=(u_{1;t},\ldots,u_{m;t})$, denote by $u_{i}({\bf u}_{t})$ the expression of $u_{i}$ as an element in $\mathbb{F}(t)_{>0}\coloneqq\mathbb{Q}_{\rm sf}(u_{1;t},\ldots,u_{m;t})$. The map $\varphi_{t}\colon u_{i}\mapsto u_{i}({\bf u}_{t})$ induces an isomorphism of semifields from $\mathbb{F}_{>0}$ to $\mathbb{F}(t)_{>0}$.
@@
    ${\bf q}^{t}(\beta)\coloneqq(\beta(u_{1;t}),\ldots,\beta(u_{m;t}))^{T}\in\mathbb{Z}^{m}.$
@@
Keep the setting in the above proposition. We call the column vector
@@
  • If we work on a $Y$-pattern $\mathcal{S}_{Y}=\{({\bf y}_{t},\widehat{B}_{t})\mid t\in\mathbb{T}_{n}\}$, we take $\mathbb{F}_{>0}$ to be the universal semifield generated by the initial $y$-variables $y_{1;t_{0}},\ldots,y_{m;t_{0}}$. Then $\mathcal{C}\coloneqq\{{\bf y}_{t}\mid t\in\mathbb{T}_{n}\}$ is a collection of charts in $\mathbb{F}_{>0}$. We usually denote by
@@
  • If we work on a cluster pattern $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$, we take $\mathbb{F}_{>0}$ to be the universal semifield generated by the initial cluster variables $x_{1;t_{0}},\ldots,x_{m;t_{0}}$. Then $\mathcal{C}\coloneqq\{{\bf x}_{t}\mid t\in\mathbb{T}_{n}\}$ is a collection of charts in $\mathbb{F}_{>0}$. We usually denote by
@@
    A tropical point ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ associated to the $Y$-pattern $\mathcal{S}_{Y}$ is an assignment of a (column) vector ${\bf g}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ in $\mathbb{Z}^{m}$ to each vertex $t$ of $\mathbb{T}_{n}$ such that
@@
    A tropical point ${\bf a}]=\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ associated to the cluster pattern $\mathcal{S}_{X}$ is an assignment of a (column) vector ${\bf a}^{t}=(a_{1}^{t},\ldots,a_{m}^{t})^{T}$ in $\mathbb{Z}^{m}$ to each vertex $t$ of $\mathbb{T}_{n}$ such that
@@
    Each tropical point ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ corresponds to a unique semifield homomorphism $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ such that
@@
    Each tropical point ${\bf a}]=\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ corresponds to a unique semifield homomorphism $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ such that